OpenStudy (anonymous):

A fireworks company has two types of rockets called Zinger 1 and Zinger 2. The polynomial -16t2+150t gives the height in feet of Zinger 1 at t seconds after launch. The polynomial: -16t2 + 165t gives the height of Zinger 2 at t seconds after launch. If the rockets are launched at the same time and both explode 6 seconds after launch, how much higher is Zinger 2 than Zinger 1 when they explode? Explain your answer to me so I could understand.

OpenStudy (anonymous):

A. 414 ft B. 990 ft C. 90 ft D. 324 ft

OpenStudy (hba):

#4. Plug t = 6 into the two formulas. Zinger 1: -16 * t^2 + 150 * t = -16 * 6^2 + 150 * 6 = Zinger 2: -16 * t^2 + 165 * t = -16 * 6^2 + 165 * 6 =

OpenStudy (anonymous):

okay 414 ?

OpenStudy (hba):

did u calculate it

OpenStudy (anonymous):

yes I got that tho ... I had used that formula but i dnt know if my answer is right.

OpenStudy (hba):

wait let me calculate

OpenStudy (anonymous):

never mind check this one distance that Z1 travels = -16(6)^2+150(6) distance that Z2 travels = -16(6)^2 + 165(6) distance between them = -16(6)^2 + 165(6) - (-16(6)^2+150(6)) = 90

OpenStudy (anonymous):

I'm i right

OpenStudy (hba):

414 is correct

OpenStudy (anonymous):

are you sure I did it again... distance that Z1 travels = -16(6)^2+150(6) distance that Z2 travels = -16(6)^2 + 165(6) distance between them = -16(6)^2 + 165(6) - (-16(6)^2+150(6)) = 90

OpenStudy (hba):

Substituting the value t = 6 in the first equation gives: (-16 * 36) + (165 * 6) = -576 + 990 = 414

OpenStudy (hba):

this is what the first polynomial gives

OpenStudy (anonymous):

ohhh okay so then my work was wrong...okay then I get it ...

OpenStudy (anonymous):

thank you again..

OpenStudy (hba):

Ur welcome